# Life Finds A Way: Emergence of Cooperative Structures in Adaptive Threshold Networks

### v2

@@ Abstract
There has been a long debate on how new levels of organization have evolved [20]. It might seem unlikely, as cooperation must prevail over competition [22]. One well-studied example is the emergence of autocatalytic sets [9], which seem to be a prerequisite for the evolution of life. Using a simple model, we investigate how varying bias toward cooperation versus antagonism shapes network dynamics, revealing that higher-order organization emerges even amid pervasive antagonistic interactions. In general, we observe that a quantitative increase in the number of elements in a system leads to a qualitative transition.
@@
Our findings extend classical random threshold and Erdős-Rényi models [8], offering new insights into adaptive systems in biological and economic contexts, with emphasis on the application to Collective Affordance Sets [18]. This framework should also be useful for making predictions that will be tested by ongoing experiments of microbial communities in soil.
@@ 1 Introduction
Microbial communities illustrate this phenomenon particularly well. Despite (or, possibly due to) immense biochemical and ecological diversity, stable interdependent communities arise repeatedly. Even when initial interactions are predominantly antagonistic, microbes find ways to coexist or even benefit each other, forming what can be viewed as "collective affordance sets" [18]: jury-rigged, self-organized structures that exploit combinations of causal properties among community members.
To empirically examine this phenomenon, ongoing experimental work aims to mix 56 bacterial and 56 fungal species, to observe if and how novel stable ecosystems emerge from a huge potential space of interaction patterns. While our primary interest is in complex soil communities of microorganisms, this concept extends beyond biology. Human economies share remarkably similar dynamics. Technological innovation and market diversity are driven by recombinations of existing products, services, and ideas. As more firms or inventors enter the market, new products and services find "niches," creating a self-reinforcing and evolving web of interdependence.
In this paper, we present an adaptive threshold network model with simple assumptions, designed to explore the conditions that foster stable, cooperative structures. Rather than focus on specific evolutionary games or particular ecological scenarios, our model investigates the broader space of possible interaction patterns, identifying conditions under which large, strongly connected communities (SCCs) reliably form and persist. Moreover, because “alternative community states can arise as a consequence of system dynamics without being driven by environmental differences" [11], the broad, environment-agnostic design of our model is well justified.
@@ 2.1 Description of Vanilla Model
  • Lifespan ($L_{S}$). Minimum number of steps a node will survive before it is culled for (i) having no in-edges or (ii) a non-positive net balance of help minus harm.
@@
  • Node traits. Each new node draws four independent traits: $h_{\text{in}},h_{\text{out}}$ or "help in" and "help out", and $\gamma_{\text{in}},\gamma_{\text{out}}$ or "harm in" and "harm out."
@@
$p_{\max}$
We set an upper‐bound parameter $p_{\max}$ (starting at $p_{\max}=0.001$ but varies in some simulations) that represents the largest possible difference in pairwise “catalytic potential” between any two nodes. For each ordered pair $(i,j)$ we draw an integer $k_{ij}\sim\mathrm{Poisson}(\lambda)$ and let the effective probability that a directed edge $i\!\to\!j$ appears be
@@
In ecological terms $k_{ij}$ could be read as the (random) count of events that would enable $i$ to facilitate or provide affordances to $j$ within the current time‐step, while $p_{\max}$ translates that count into a probability of edge-formation.
@@ 3 Experiments & Results
In a systematic way, we explore how shifting the parameters (lifespan, harm-to-help ratio, bind-range) influence the emergence and persistence of strongly connected components (SCCs) in our directed network. Our primary interest was the evolution and robustness of cooperative structures (large SCCs) in networks characterized by varying degrees of antagonism (harm-to-help ratio). Eigen Analysis was performed to give a higher level characterization of the network’s behavior short and long run behavior.
@@ 3.1 Harm-To-Help Ratio and Lifespan Dynamics
3.1.1 Below 0.5 Harm-To-Help Ratio
When cooperative interactions dominate (harm ratio $\rho<0.5$), we observed exponential growth in directed edge formation, resulting in a linearly increasing SCC size relative to the number of nodes in the network. This scenario characterizes a cooperative and expansive community where nodes not only persist, but integrate rapidly into a collective structure. Once such a community emerges, with a majority of help edges, it is would be difficult for its elements to be removed.
Figures 3,4,5 show the results of the simulation from solely increasing the Harm-To-Help ratio from $\rho=0.3,0.5,0.7$ with all other parameters fixed. At first, we can restrict our view to the differences in the number of nodes and edges at step $t$, the largest SCC size and largest SCC fraction.
@@
3.1.2 When Harm-To-Help >0.5>0.5
For harm-to-help ratios exceeding 0.5, a strongly connected community still emerges and persists, provided nodes’ lifespan surpasses $\approx 10$ timesteps. Below this number, an SCC will fail to form reliably if $\rho>0.7$.
@@ 3.2 Influence of Binding Chance
Adjusting the binding chance, $p_{max}$ which scales the likelihood threshold for interactions, primarily impacts the rate of convergence toward the final dynamic state. Increasing this parameter leads more quickly to equilibrium behaviors without substantially altering the qualitative behavior of SCC growth. However, in scenarios where the harm ratio exceeded 0.5, a higher binding chance increased the fraction of nodes in the SCC, but overall slightly decreased the number of nodes in the equilibrium size of the SCC. This suggests that increasing the likelihood of interaction partially counteracts the destabilizing effect of widespread antagonism.
@@ 3.3 Node Age and Persistence
We noted a ‘demographic’ shift among highly-connected nodes (in the top 10% of in-degree) as antagonism increased. Specifically, and perhaps not surprisingly, networks with higher antagonism were increasingly dominated by younger nodes. High-harm environments shorten the node survival significantly, reducing the presence of “older" and more established nodes. Despite this shorter persistence, large SCCs remain viable and indicate a dynamically renewing structure of cooperation, which is resilient amidst high node turnover.
@@
Figure 7: Log(age) vs Log(degree) after completing 10,000 steps of the simulation. $p_{max}=0.003$. $\rho=0.2,0.5,0.8$.
3.4 Eigen Analysis of Adjacency Matrix
We first generate an adjacency matrix by flattening the MultiGraph to a weighted graph by summing the number of edges: +1 for help edges and -1 for harm edges. For example entry $A_{ij}$ will hold +3 if there are 5 helpful and 2 harmful edges from $i\to j$. Recall that at most 6 edges of one type (help or harm) can form between two nodes.
At each simulation tick we run a 20-step power iteration to approximate the leading eigenpair. A simple deflation then yields the sub-dominant eigenvalue. The resulting spectral radius, gap, and the angle between successive principal eigenvectors give a coarse but fast-to-compute fingerprint of network cohesion and regime shifts. While this heuristic skips the full, potentially complex spectrum and may miss fine detail, it is accurate enough for the qualitative trends we report and avoids the heavy cost of exact eigendecomposition on every time-step.
[…]
Summary:
  • Spectral angle = measures structural change between network states.
@@ 4.1 Robustness of Cooperation under High Antagonism
Our model demonstrates robustness of cooperative structures even when antagonistic interactions (harmful edges) significantly outnumber supportive interactions. This finding could be an extension of the concept of ‘islands of cooperation’ in [21]. Specifically, even with high harm-to-help ratios ($\rho>0.5$), SCCs consistently emerge and persist at steady-state equilibria. This implies the presence of inherent resilience within the network structure, which allows it to dynamically reconfigure to maintain cooperative viability despite continuous antagonistic pressure. The results on node “age" vs degree distribution lend another perspective on the long run dynamics, which we discuss below.
@@ 4.2 Critical Thresholds and Phase Transitions
A primary result from our simulations is the clear emergence of a phase transition at the harm-to-help ratio $\rho\approx 0.5$. Below this threshold, the strongly connected component (SCC) grows linearly and steadily with the total node population, indicating robust, widespread cooperation. Above this threshold, however, SCC growth transitions from indefinite linear expansion to a bounded steady state, characterized by persistent fluctuations. The shift marks a clear qualitative boundary in the dynamics of the system.
Furthermore, we observe another critical threshold related to the bind chance parameter ($p_{\max}$). When $p_{\max}$ is sufficiently low (around $0.002$), there emerges a clear criticality in network density: a minimum population size (approximately 200 nodes, just after 500+ steps, in our simulations, see Figure 8) must first be reached before an SCC becomes viable and begins growing rapidly. This behavior resembles classical phase transitions in percolation theory and random graph models, where a threshold density of connections is required for a "giant" connected component to emerge. [8] Explorations of these transitions using Haantjes curvature [27] have produced promising preliminary results, indicating that mean curvature may reliably increase prior to the formation of the emerging SCC.
Both phase transitions identified here highlight intriguing qualitative boundaries and merit deeper analytic investigation. Future work should include refined analytical treatments and fine-grained numerical explorations, particularly around the critical harm-to-help ratio region $0.5<\rho<0.6$, to conclusively determine the precise nature and location of these transitions.
Figure 8: $p_{max}=0.002,\rho=0.2$
@@ 4.3 Eigen Analysis
Eigen Analysis of the network’s adjacency matrix provides more insight into structural and dynamical features. The principal eigenvalue (found by power iteration in NumPy) measures overall connectivity strength along the dominant eigenvector direction, which reflects the magnitude and stability of cooperation (or antagonism) in the network. Higher eigenvalues indicate denser and strongly coupled cooperative or competitive interactions. In our simulations, rising mean eigenvalues (observed with $\rho<0.5$ ) corresponded to increasingly stable and interconnected SCCs.
@@ 4.4 Age Distribution and Community Stability
The interplay between node lifespan and network antagonism produces distinctive patterns in node age (measured in time - birth-step) distributions. Under conditions of low antagonism (low $\rho$), older nodes with high connectivity become established, reflecting stable core communities. Conversely, increased antagonism rapidly shifts this balance toward younger nodes, as frequent node removal due to harmful interactions reduces node longevity. Interestingly, despite the reduction in older, established nodes, SCCs remain viable by continuously replenishing their core through newly added nodes. This activity reflects a structurally dynamic equilibrium rather than a fixed, static composition. Investigating the details behind the bands and clusters we see in the logarithmic plot may yield more insights. It is notable that at $\rho=0.5$ we see a cluster of nodes that are persisting most of the simulation, but have a reduced overall degree.

### v3

@@ Abstract
There has been a long debate on how new levels of organization have evolved [20]. The evolution of higher-level organization can appear unlikely, since cooperation must prevail over competition [22]. One well-studied example is the emergence of autocatalytic sets [9], which are often considered a prerequisite for the evolution of life.
Using a simple model, we investigate how varying bias toward cooperation versus antagonism shapes network dynamics, revealing that higher-order organization emerges even amid pervasive antagonistic interactions. In general, we observe that a quantitative increase in the number of elements in a system leads to a qualitative transition.
@@
Our findings extend classical random threshold and Erdős-Rényi models [8], offering new insights into adaptive systems in biological and economic contexts, with emphasis on the application to Collective Affordance Sets [18]. This framework will also be useful for making predictions that will be tested by ongoing experiments of microbial communities in soil.
journal: Theoretical Biology
[…]
[label1]organization=SUNY Finger Lakes CC, addressline=3325 Marvin Sands Dr, city=Canandaigua, postcode=14424, state=NY, country=USA
[…]
[label2]organization=SUNY Binghamton, addressline=4400 Vestal Pkwy E, city=Binghamton, postcode=13902, state=NY, country=USA
[…]
[label3]organization=Institute for Systems Biology, addressline=401 Terry Ave N, city=Seattle, postcode=98109, state=WA, country=USA
@@ 1 Introduction
Microbial communities illustrate this phenomenon particularly well. Despite (or, possibly due to) immense biochemical and ecological diversity, stable interdependent communities arise repeatedly. Even when initial interactions are predominantly antagonistic, microbes find ways to coexist or even benefit each other, forming what can be viewed as ”collective affordance sets” [18]: jury-rigged, self-organized structures that exploit combinations of causal properties among community members.
To empirically examine this phenomenon, ongoing experimental work aims to mix 56 bacterial and 56 fungal species, to observe if and how novel stable ecosystems emerge from a huge potential space of interaction patterns. While our primary interest is in complex soil communities of microorganisms, this concept extends beyond biology. Human economies share remarkably similar dynamics. Technological innovation and market diversity are driven by recombinations of existing products, services, and ideas. As more firms or inventors enter the market, new products and services find ”niches,” creating a self-reinforcing and evolving web of interdependence.
Cooperative and interdependent structures in microbial communities need not arise from explicitly altruistic strategies. A prominent alternative framework is the Black Queen Hypothesis, which explains the emergence of metabolic dependencies through adaptive gene loss and the retention of “leaky” functions by a subset of community members. In this view, organisms may reduce individual costs by abandoning costly functions while relying on others to supply shared resources, producing cooperation-like patterns without cooperative intent. The present work is not intended to model gene loss or metabolic specialization directly. Rather, it focuses on the dynamical consequences of structured help and harm interactions at the network level, abstracting away from the evolutionary pathways by which such interactions arise. As such, the framework developed here should be viewed as complementary to Black Queen–type mechanisms, addressing how interdependence (once present) shapes the stability and organization of communities.
In this paper, we present an adaptive threshold network model with simple assumptions, designed to explore the conditions that foster stable, cooperative structures. Rather than focus on specific evolutionary games or particular ecological scenarios, our model investigates the broader space of possible interaction patterns, identifying conditions under which large, strongly connected communities (SCCs) reliably form and persist. Moreover, because “alternative community states can arise as a consequence of system dynamics without being driven by environmental differences” [11], the broad, environment-agnostic design of our model is well justified.
@@ 2.1 Description of Vanilla Model
  • 2.
    Lifespan ($L_{S}$). Maximum number of steps a node may survive before it is culled for (i) having no in-edges or (ii) a non-positive net balance of help minus harm.
  • 3.
@@
  • 5.
    Node traits. Each new node draws four independent traits: $h_{\text{in}},h_{\text{out}}$ or ”help in” and ”help out”, and $\gamma_{\text{in}},\gamma_{\text{out}}$ or ”harm in” and ”harm out.”
@@
2.1.1 $p_{\max}$ parameter
We set an upper‐bound parameter $p_{\max}$ (starting at $p_{\max}=0.001$ but varies in some simulations) that represents the largest possible difference in pairwise “catalytic potential” between any two nodes. For pair-specific variability, at the time a node $i$ is added, each ordered pair of nodes $(i,j)$ we draw an integer $k_{ij}\sim\mathrm{Poisson}(\lambda)$ and let the effective probability that a directed edge $i\!\to\!j$ and/or $j\!\to\!i$ appears be
@@ 2.1.1 $p_{\max}$ parameter
To clarify, this $P_{ij}$ is set once between each entering node and every existing node. In ecological terms $k_{ij}$ could be read as the (random) count of events that would enable $i$ to facilitate or provide affordances to $j$ within the current time‐step, while $p_{\max}$ translates that count into a probability of edge-formation.
2.2 Eigen Analysis of Adjacency Matrix
At each simulation tick, we generate an adjacency matrix by flattening the MultiGraph to a weighted graph by summing the number of edges: +1 for help edges and -1 for harm edges. For example entry $A_{ij}$ will hold +3 if there are 5 helpful and 2 harmful edges from $i\to j$. Recall that at most 6 edges of one type (help or harm) can form between two nodes. We then approximate the dominant eigenpair (eigenvector and corresponding eigenvalue) $(\hat{\lambda}_{1},\hat{v}_{1})$ using a fixed-iteration power method with normalization at each step and a Rayleigh-quotient estimate,
$\hat{\lambda}_{1}\;=\;\frac{\hat{v}_{1}^{\top}A_{t}\hat{v}_{1}}{\hat{v}_{1}^{\top}\hat{v}_{1}},$
terminating after at most 20 iterations or earlier if successive Rayleigh estimates satisfy $|\hat{\lambda}_{1}^{(k)}-\hat{\lambda}_{1}^{(k-1)}|<10^{-6}$.
To obtain a computationally efficient approximation of the sub-dominant eigenvalue, we perform a rank-one deflation by subtracting the estimated dominant mode,
$B_{t}\;=\;A_{t}\;-\;\hat{\lambda}_{1}\,\hat{v}_{1}\hat{v}_{1}^{\top},$
and reapply the same iterative procedure to $B_{t}$ to estimate $\hat{\lambda}_{2}$. We then report the spectral gap as $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$ and track the principal-angle change between successive dominant eigenvectors as a proxy for turnover in the network core. Because $A_{t}$ is generally non-symmetric, these spectral quantities should be interpreted as efficient heuristics for cohesion and regime change rather than as exact spectral characterizations at each time step.
[…]
Summary:
  • 1.
  • 2.
    Spectral angle = measures structural change between network states.
  • 3.
@@ 3 Experiments & Results
In a systematic way, we explore how shifting the parameters (lifespan, harm-to-help ratio, binding chance) influence the emergence and persistence of strongly connected components (SCCs) in our directed network. Our primary interest was the evolution and robustness of cooperative structures (large SCCs) in networks characterized by varying degrees of antagonism (harm-to-help ratio). Eigen Analysis was performed to give a higher level characterization of the network’s behavior short and long run behavior.
@@ 3.1 Harm-To-Help Ratio and Lifespan Dynamics
3.1.1 Low antagonism ($\rho$ below 0.5~0.5)
When cooperative interactions dominate (harm ratio $\rho<0.5$), we observed exponential growth in directed edge formation, resulting in a linearly increasing SCC size relative to the number of nodes in the network. This scenario characterizes a cooperative and expansive community where nodes not only persist, but integrate rapidly into a collective structure. Once such a community emerges, with a majority of help edges, it would be difficult for its elements to be removed.
Figures 4,5,6 show the results of the simulation from solely increasing the Harm-To-Help ratio from $\rho=0.3,0.5,0.7$ with all other parameters fixed. At first, we can restrict our view to the differences in the number of nodes and edges at step $t$, the largest SCC size and largest SCC fraction.
3.1.2 High antagonism ($\rho$ above 0.5~0.5)
For harm-to-help ratios exceeding 0.5, a strongly connected community still emerges and persists, provided nodes’ lifespan surpasses $\approx 10$ timesteps. Below this number, an SCC will fail to form reliably if $\rho$ is sufficiently high. See Figure 3.
Figure 3: Above harm ratio $\rho>0.5$, lifespan below $L=10$ will not reliably produce an SCC.
@@ 3.2 Influence of Binding Chance
Adjusting the binding chance, $p_{max}$ which scales the likelihood threshold for interactions, primarily impacts the rate of convergence toward the final dynamic state (Figure 11). Increasing this parameter leads more quickly to equilibrium behaviors without substantially altering the qualitative behavior of SCC growth.
However, in scenarios where the harm ratio exceeded 0.5, a higher binding chance increased the fraction of nodes in the SCC, but overall slightly decreased the number of nodes in the equilibrium size of the SCC. This suggests that increasing the likelihood of interaction partially counteracts the destabilizing effect of widespread antagonism. See Figure 7.
@@ 3.3 Node Age and Persistence
We noted a ‘demographic’ shift among highly-connected nodes (in the top 10% of in-degree) as antagonism increased. Specifically, and perhaps not surprisingly, networks with higher antagonism were increasingly dominated by younger nodes. High-harm environments shorten the node survival significantly, reducing the presence of “older” and more established nodes. Despite this shorter persistence, large SCCs remain viable and indicate a dynamically renewing structure of cooperation, which is resilient amidst high node turnover.
@@
Figure 8: Log(age) vs Log(degree) after completing 10,000 steps of the simulation. Each point represents a node in the network at N=10,000 where $p_{max}=0.003$, at varying levels of harm: $\rho=0.2,0.5,0.8$. See video animation link for time-lapse visualization.
@@ 4 Discussion
Clarifying Scope of Vanilla Model
The framework introduced here operates deliberately at a higher level of abstraction. Rather than modeling specific biochemical, ecological, or evolutionary mechanisms that generate interactions (e.g., metabolic exchange, spatial structuring, gene loss, or adaptive specialization), the model treats the interaction network itself as the primary object of study. In this sense, it explores the space of possible help–harm configurations that such mechanisms may produce, rather than the mechanisms themselves. This abstraction allows us to identify network-level regularities—such as the persistence of large strongly connected components under mixed antagonism and support—that are potentially shared across microbial, economic, and other complex systems. At the same time, this perspective necessarily leaves unresolved questions about how particular interaction patterns arise, how they are modulated by spatial or metabolic constraints, or how they evolve over long timescales. The present results should therefore be interpreted as constraints on, rather than substitutes for, more detailed mechanistic models: any system whose interaction structure falls within this class would be expected to exhibit similar collective behavior, regardless of the underlying generative process.
@@ 4.1 Robustness of Cooperation under High Antagonism
Our model demonstrates robustness of cooperative structures (large SCCs) even when antagonistic interactions (harmful edges) significantly outnumber supportive interactions. This finding could be an extension of the concept of ‘islands of cooperation’ in [21]. Specifically, even with high harm-to-help ratios ($\rho>0.5$), SCCs consistently emerge and persist at steady-state equilibria. This implies the presence of inherent resilience within the network structure, which allows it to dynamically reconfigure to maintain cooperative viability despite continuous antagonistic pressure. The results on node “age” vs degree distribution lend another perspective on the long run dynamics, which we discuss later.
The interaction structure considered here assumes that benefits and harms are effectively public, an assumption that simplifies analysis but departs from many microbial systems in which goods are partially privatized or retained to varying degrees. In natural communities, metabolites may be incompletely shared, spatially constrained, or dynamically regulated, producing interaction strengths that vary over time and context.
@@ 4.2 Critical Thresholds and Phase Transitions
A central outcome of our simulations is the emergence of a phase transition in the system behavior at the harm-to-help ratio $\rho\approx 0.6$. (Figure 9, 10) For values of $\rho$ below this threshold, the strongly connected component (SCC) grows steadily with the total node population, reflecting a regime of sustained interaction and widespread mutual support. Above this threshold, however, SCC growth transitions from indefinite expansion to a bounded, fluctuating steady state.
This transition is evident in the late-time population growth rate (measured over the final 500 steps), which decreases sharply and crosses zero as a function of $\rho$ (Figure 9). A corresponding collapse in the mean network degree is shown in Figure 10, indicating a loss of interaction density that accompanies the change in growth dynamics. Together, these results identify a clear qualitative boundary separating the regimes of population growth from bounded population dynamics.
Figure 9: The late-time (last 500 steps) population growth rate $g$ as a function of the harm-to-help ratio $\rho$. For small $\rho$, the system exhibits sustained positive growth. As $\rho$ increases, the growth rate decreases sharply and crosses zero near $\rho\approx 0.6$, marking a transition from sustained growth to a bounded population regime.
Figure 10: Mean total degree (in + out) as a function of the harm-to-help ratio $\rho$. As $\rho$ increases, the average connectivity of the network decreases sharply, with a pronounced collapse occurring near $\rho\approx 0.6$. This reduction in interaction density coincides with the transition from sustained to bounded population growth observer in the above figure, suggesting the loss of sufficient interaction capacity underlies the change in population dynamics.
While the harm-to-help ratio $\rho$ determines whether the system exhibits sustained or bounded growth, the binding chance $p_{max}$ governs the scale at which global coordination emerges. Figure 11 shows the fraction of nodes in the largest strongly connected component as a function of population size for different values of $p_{max}$. Increasing $p_{max}$ sharply reduces the critical population size required for SCC formation, indicating a transition in coordination onset rather than final network structure. This behavior resembles classical phase transitions in percolation theory and random graph models, where a threshold density of connections is required for a ”giant” connected component to emerge. [8] Together, these results demonstrate that $\rho$ and $p_{max}$ control distinct but complementary aspects of the system’s macroscopic behavior.
Figure 11: Onset of the largest strongly connected component (SCC) as a function of population size for varying bind chance $p_{max}$ ($\rho=0.2$, $L_{S}=100$). Increasing $p_{max}$ shifts the emergence of global connectivity to progressively smaller population sizes, indicating a transition in the onset of coordination. All curves converge to a near-unity SCC fraction, showing that $p_{max}$ primarily controls the timing of collective organization.
@@ 4.3 Eigen Analysis
Eigen Analysis of the network’s adjacency matrix provided us more insight into structural and dynamical features. The principal eigenvalue (approximated by method outlined in Section 2.2) measures overall connectivity strength along the dominant eigenvector direction, which reflects the magnitude and stability of cooperation (or antagonism) in the network. Higher eigenvalues indicate denser and strongly coupled cooperative or competitive interactions. In our simulations, rising mean eigenvalues (observed with $\rho<0.5$ ) corresponded to increasingly stable and interconnected SCCs.
@@ 4.4 Age Distribution and Community Stability
The interplay between node lifespan and network antagonism produces distinctive patterns in node age (measured in time - birth-step) distributions (Figure 8). Under conditions of low antagonism (low $\rho$), older nodes with high connectivity become established, reflecting stable core communities. Conversely, increased antagonism rapidly shifts this balance toward younger nodes, as frequent node removal due to harmful interactions reduces node longevity. Interestingly, despite the reduction in older, established nodes, SCCs remain viable by continuously replenishing their core through newly added nodes. This activity reflects a structurally dynamic equilibrium rather than a fixed, static composition. Investigating the details behind the bands and clusters we see in the logarithmic plot may yield more insights. It is notable that at $\rho=0.5$ we see a cluster of nodes that are persisting most of the simulation, but have a reduced overall degree.